OpenStudy (is3535):

HELP HELP HELPWhat is the value of the expression 3/3^-2

OpenStudy (is3535):

@Trisaba

OpenStudy (is3535):



OpenStudy (is3535):

@KendrickLamar2014

OpenStudy (anonymous):

3^3

OpenStudy (kendricklamar2014):

What have you done so far?

OpenStudy (is3535):

3 * 3

OpenStudy (kendricklamar2014):

Here are the steps: \[\left(\begin{matrix}3^1 \\ 3^-2\end{matrix}\right)\] \[3^1(-2)\]

OpenStudy (kendricklamar2014):

\[= 3^3\] \[Answer= 27\]

OpenStudy (kendricklamar2014):

Do you understand @is3535 ?

OpenStudy (is3535):

yea

OpenStudy (kendricklamar2014):

Glad I could help!

OpenStudy (is3535):

ok